Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
The sum of two consecutive even integers is 86. Find the larger number.

Answer by oberobic(2304)

The 5-Step Method
1) Let x (or some other letter) represent the unknown quantity.
2) Translate the English to mathematics and form an equation.
3) Solve this equation.
4) Check this result by substituting it into the original statement of the problem.
5) Write a conclusion.


Step 1.
x = an even integer
y = an even integer that is consecutive to x
It is better to say: y = x+2


Step 2.
The sum of two consecutive even integers is 86.
x+y = 86


Step 3.
substituting y=x+2
x + x+2 = 86
2x+2 = 86
2x = 84
x = 42
y = x+2 = 42+2= 44


Step 4.
checking the solution
x + y = 86?
42 + 44 = 86
correct.


Step 5.
Reviewing the question, it asks "what is the larger number?"
So, our conclusion should answer the question:
The larger of two consecutive even integers that sum to 86 is 44.


Done.